Amazon Beverages produces and bottles a line of soft drinks using exotic fruits from Latin America and Asia. The manufacturing process entails mixing and adding juices and coloring ingredients at the bottling plant, which is a part of Mixing Division. The finished product is packaged in a company-produced glass bottle and packed in cases of 24 bottles each.

 

Because the appearance of the bottle heavily influences sales volume, Amazon developed a unique bottle production process at the company’s container plant, which is a part of Container Division. Mixing Division uses all of the container plant’s production. Each division (Mixing and Container) is considered a separate profit center and evaluated as such. As the new corporate controller, you are responsible for determining the proper transfer price to use for the bottles produced for Mixing Division.

 

At your request, Container Division’s general manager asked other bottle manufacturers to quote a price for the number and sizes demanded by Mixing Division. These competitive prices follow.

 

VolumeTotal PricePrice per Case
440,000 equivalent casesa$3,344,000
$7.60
880,000
5,808,000

6.60
1,320,000
7,656,000

5.80

 

a An equivalent case represents 24 bottles.

 

Container Division's cost analysis indicates that it can produce bottles at these costs.

 

VolumeTotal CostCost per Case
440,000 equivalent cases$2,776,000
$6.31
880,000
4,712,000

5.35
1,320,000
6,648,000

5.04

 

These costs include fixed costs of $840,000 and variable costs of $4.40 per equivalent case. These data have caused considerable corporate discussion as to the proper price to use in the transfer of bottles from Container Division to Mixing Division. This interest is heightened because a significant portion of a division manager’s income is an incentive bonus based on profit center results.

 

Mixing Division has the following costs in addition to the bottle costs.

 

VolumeTotal CostCost per Case
440,000 equivalent cases$1,840,000
$4.18
880,000
2,640,000

3.00
1,320,000
3,440,000

2.61

 

The corporate marketing group has furnished the following price–demand relationship for the finished product:

 

Sales VolumeTotal Sales
Revenue
Sales Price
per Case
440,000 equivalent cases$8,976,000
$20.40
880,000
16,192,000

18.40
1,320,000
20,328,000

15.40

 

Required:

a. Amazon Beverages has used market price–based transfer prices in the past. Using the current market prices and costs and assuming a volume of 1.32 million cases. Calculate operating profits for Container Division, Mixing Division, Amazon Beverages.

b-1. Calculate operating profits for Container, Mixing and Amazon Beverages for volumes of 440,000, 880,000 and 1,320,000 cases.

b-2. Which volume of production is the most profitable for Container, Mixing and Amazon Beverages?
Answer #1

A) 7,656- 6,648= 1,008 profit for container

7,656+ 3,440= 11,096    20,328- 11,096= 9,232 profit for mixing division

6,648+ 3,440= 10,088    20,328-10,088= 10,240 profit for Amazon division


                            440,000 cases        880,000 cases          1,320,000 cases

Container profit           568                        1,096                         1,008

Mixing profit               $3,792                    $7,744                     $9,232

Amazon Beverage profit$4,360               $8,840                       $10,240


container profit

440: 3,344- 2,776= 568

880: 5,808- 4,712= 1,096

1,320: 7,656- 6,648= 1,008


mixing division

step 1) find cost

1,840+ 3,344= 5,184

2,640+ 5,808= 8,448

3,440+ 7,656= 11,096


440: 8,976(from corporation chart)- 5,184= 3,792

880: 16,192- 8,448= 7,744

1,320: 20,328- 11,096= 9,232


corporation division

1,840+ 2,776= 4,616

4,712+ 2,640= 7,352

3,440+ 6,648= 10,088


440: 8,976- 4,616= 4,360

880: 16,192- 7,352= 8,840

1,320: 20,328- 10,088= 10,240


B2) most profitable volume for container: 880,000

most profitable volume for mixing  1,320,000

most profitable volume for Amazon  1,320,000